OpenStudy (anonymous):

Given the function f(x) = -2x + 8, find x if f(x) = 22. Answer -36 -15 -7 52

hero (hero):

f(x) = -2x + 8 22 = -2x + 8 solve for x

OpenStudy (anonymous):

dont get it

OpenStudy (anonymous):

f(x) = -2x + 8 = 22 -> -2x = 22 - 8 => x = 14/ -2 = -7

hero (hero):

You don't get that f(x) = 22 or you don't know how to solve for x when given: 22 = -2x + 8 ?